If a length of a cylindrical conductor increased by 20% keeping the volume constant.then percentage change in resistance of the conductor is 11×n% the value of n 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
as volume is constant
 
V1 = V2
 
A1 * L1 = A2 * L2
 
L2 / L1 = A 1 / A2 = 1.02
 
Now
 
R = rho * L /A
 
hence
 
R1 / R2 = L1 *A2/ L2 *A1
 
R1 / R2 = 1 / 1.02^2
 
R1 / R2 = 1 / 1.0404
 
R2 = 1.0404R1